OpenStudy (anonymous):

Find an equation of tangent line at indicated point on graph y=f(x)=4 sqrt(x) -x+ 7 (x,y=16,7)

OpenStudy (anonymous):

\[f(x)=4\sqrt{x}-x+7\] \[f'(x)=\frac{2}{\sqrt{x}}-1\] find \(f'(16)\) for the slope, then use the "point-slope" formula for the equation

OpenStudy (anonymous):

the slope should come out to have 2 values 0.5 and -1.5 but how come the solution is only -1.5 slope. Do you know why